The beam of parallel light rays falls onto the surface of water under the angle



60°



. The width of the beam in the air is 10 cm. Find the width of the beam in the water.

1
Expert's answer
2022-01-12T08:31:49-0500

"n=\frac{\sin\alpha}{\sin\beta},\implies \beta=40.5°,"

"b=\frac{a\cos\beta}{\cos\alpha}=15.2~cm."